40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 16 02:11:40 2014, Jenne, Update, LSC, CARM and DARM on IR signals, boosts engaged 
    Reply  Wed Apr 16 02:29:30 2014, ericq, Update, LSC, CARM and DARM on IR signals, boosts engaged carmOLTF.pdfboostPlot.png
       Reply  Thu Apr 17 00:49:06 2014, Jenne, Update, LSC, CARM and DARM on IR signals, boosts engaged Lockloss_DARMgainTooHigh_119am.pngLockloss_DARMgainTooHigh_127am.pngSmoothCARMandDARMtransitions_LSCboosts.png
Message ID: 9819     Entry time: Thu Apr 17 00:49:06 2014     In reply to: 9818
Author: Jenne 
Type: Update 
Category: LSC 
Subject: CARM and DARM on IR signals, boosts engaged 

I looked at 2 of the locklosses from last night, (1:19am and 1:27am), and saw that for both, the DARM loop started to oscillate just before we lost lock.  In the trials tonight, we were more watchful of gain peaking.

Here is the 1:19am lockloss

Lockloss_DARMgainTooHigh_119am.png

And here is the 1:27am lockloss

Lockloss_DARMgainTooHigh_127am.png


 So you can see what we were doing, and what the effect was, here is a few minutes of data just before the 1:27am lockloss. The times I note below are rough, but should give you an idea of what happened at which time.

0 sec:  Arms are held on resonance with ALS.

10 sec:  CARM offset of 3nm added.

20 sec:  PRM restored, one flash, then PRMI acquires lock.

30 sec:  CARM offset reduced to 2nm, transmitted powers are about 0.1

50 sec:  Transition CARM to 1/sqrt(trans) signals.  Note that we are using the high gain Thorlabs PD here, so the noise is better than last Thursday.

60-110 sec:  CARM offset reduction to about 1nm.

110 sec:  CARM's LSC low frequency boost engaged.

120 sec:  DARM transitioned to AS55Q.

170 sec:  DARM's LSC low frequency boost engaged.

SmoothCARMandDARMtransitions_LSCboosts.png

ELOG V3.1.3-